OpenStudy (kellyspeakslouder):

f(x) = 3x + 2; g(x) = 3x - 5 Find f/g. A.(f/g)(x) = Quantity three x minus five divided by three x plus two.; domain {x|x ≠ - Two over three. } B. (f/g)(x) = Quantity three x plus two divided by three x minus five.; domain {x|x ≠ Five over three. } C. (f/g)(x) = Quantity three x minus five divided by three x plus two; domain {x|x ≠ Five over three. } D. (f/g)(x) = Quantity three x plus two divided by three x minus five; domain {x|x ≠ - Two over three. }

satellite73 (satellite73):

yikes

OpenStudy (kellyspeakslouder):

for real.. it's a headache

satellite73 (satellite73):

equation editor would help a great deal, but put one over the other

satellite73 (satellite73):

like \[\frac{3x+2}{3x-5}\] that is all

OpenStudy (shadowlegendx):

lol xD

satellite73 (satellite73):

oh i see you have another part, the "domain"

satellite73 (satellite73):

to find it, set \[3x-5=0\] and solve for \(x\)in two steps only then say "all real numbers except that one"

satellite73 (satellite73):

let me know when you get it

OpenStudy (kellyspeakslouder):

im sorry, I don't get it I'm really confused

satellite73 (satellite73):

ok do you get the \[\frac{f}{g}\] part?

OpenStudy (kellyspeakslouder):

yes

satellite73 (satellite73):

ok so that is taken care of, good now there is one more part, the "domain" lets take it slow

satellite73 (satellite73):

you cannot "divide by zero" (i tired it once, had a headache for a week) which is the math teacher way of saying that the denominator of your fraction \(3x-5\) cannot be equal to zero in practical terms what this means is that you have to solve the equation \[3x-5=0\] for \(x\) do you know how to do that?

satellite73 (satellite73):

"no" is a fine answer, just asking

OpenStudy (kellyspeakslouder):

no..

satellite73 (satellite73):

ok lets do it

satellite73 (satellite73):

\[3x-5=0\] add 5 to both sides get \[3x=5\] then divide both sides by 3 to finish with \[x=\frac{5}{3}\]

satellite73 (satellite73):

like i said, two steps only add 5, divide by 3 so your domain is all real numbers except \(x=\frac{5}{3}\)

OpenStudy (kellyspeakslouder):

ohhh okay

OpenStudy (kellyspeakslouder):

so it would be C?

OpenStudy (kellyspeakslouder):

@zepdrix

OpenStudy (kellyspeakslouder):

hello?